OpenStudy (anonymous):

A building measures 50 feet in length and w feet in width. Which of the following is an expression for the perimeter of the building?

OpenStudy (anonymous):

2(50+w) = P

OpenStudy (angela210793):

P=2l+2w=2*50+2l=100+2l